### v1

# Large Language Models as Proxies for Theories of Human Linguistic Cognition

Abstract
We consider the possible role of current large language models (LLMs) in the study of human linguistic cognition. We focus on the use of such models as proxies for theories of cognition that are relatively linguistically-neutral in their representations and learning but differ from current LLMs in key ways. We illustrate this potential use of LLMs as proxies for theories of cognition in the context of two kinds of questions: (a) whether the target theory accounts for the acquisition of a given pattern from a given corpus; and (b) whether the target theory makes a given typologically-attested pattern easier to acquire than another, typologically-unattested pattern. For each of the two questions we show, building on recent literature, how current LLMs can potentially be of help, but we note that at present this help is quite limited.
Keywords: Learning, Poverty of the stimulus, Linguistic typology, Large language models
1 Introduction
The question of whether current large language models (LLMs) might be relevant scientifically for the study of human linguistic cognition (HLC) is often framed in terms of a perceived threat that LLMs pose to generative linguistics. On the one hand, some researchers have suggested that these models discredit prevailing views within linguistics (Piantadosi, 2023), while others have suggested that LLMs can teach us nothing about HLC (Chomsky et al., 2023, Moro et al., 2023a).
Here we chart a more collaborative path forward for the integration of LLMs into the study of HLC. While the idea that current LLMs are themselves good theories of HLC is untenable (as reviewed below), we note that LLMs are a potentially useful tool that can inform the evaluation of theories of HLC. In this capacity, LLMs can serve as proxies for explicit theories of HLC that are different from LLMs in important ways but are still relatively unbiased with respect to linguistic representations and learning. We take it that much of the recent literature on the potential cognitive relevance of LLMs takes current models to support relatively linguistically neutral theories of this kind. To the extent that this use of LLMs as proxies is justified, they can assist in comparing such relatively unbiased theories of HLC with more traditional theories of HLC coming out of theoretical linguistics, where the combination of representations and learning is often highly biased in favor of linguistic patterns.
We start in section 2 by briefly reviewing some considerations pertaining to the role of LLMs in studying HLC. We start with a reminder that the study of HLC, like the rest of empirical science, proceeds through inference to the best explanation, with explicit theories competing in light of known observations. We review the recent idea that LLMs themselves can serve as better explanations than generative theories and then summarize some of the reasons noted in the literature to reject this idea. We then consider the idea mentioned above that the best explanation is that HLC is quite different from current LLMs but still involves representations and learning that are relatively linguistically neutral. It is the potential of LLMs to serve as proxies for such a relatively linguistically-neutral theory that will preoccupy us in the remainder of the paper, and in particular the question of whether evidence from LLMs can bear on the question of whether the linguistically-neutral theory is a better explanation than generative approaches. In section 3 we ask – building on Wilcox et al. 2023 and Lan et al. 2024b, among others – whether evidence from LLMs suggests that a relatively linguistically-neutral theory of HLC might succeed in acquiring various linguistic patterns from a developmentally-realistic corpus. In section 4 we ask – building on Kallini et al. 2024 and others – whether evidence from LLMs suggests that a relatively linguistically-neutral theory of HLC makes attested languages easier to acquire than various unattested variants of those languages. In both cases, we will see that the evidence from LLMs gives no reason to think that a linguistically-neutral theory of HLC could account for the data, which in turn means that current LLMs contribute little to the ability of such linguistically-neutral theories to challenge generative theories.
2 Background: LLMs, HLC, and the inference to the best explanation
Being part of empirical science, the study of HLC proceeds through inference to the best explanation: explicit theories are evaluated in light of the observations, and those theories that explain the data best are preferred to the other contenders. The theories coming out of generative linguistics over the past 70 years or so are routinely evaluated in this way. These theories are often strongly linguistically biased, in the sense that they make it relatively easy to represent and learn knowledge of the kind found across languages and hard or impossible to represent or learn many other patterns. Let us use $H_{1}$ to refer to some concrete instantiation of the generative approach. We will return to the biases embodied by the architecture and representations of $H_{1}$ in the following sections.
More recently, Piantadosi (2023) has suggested that current LLMs should be treated as explicit theories of HLC that should be taken seriously as part of this evaluation. Following Fox and Katzir (2024) we refer to this new approach to HLC as the LLM Theory, and we discuss it in Section 2.1, through a concrete instantiation we will call $H_{2}$. Perhaps in light of the glaring inadequacies of the LLM Theory, some recent literature, including Wilcox et al. 2023 and Mahowald et al. 2024, has advocated for a revision of generative linguistics on the basis of LLMs but without committing to the LLM Theory. We refer to this idea as the Proxy View:
[…]
Since the literature that promotes the Proxy View generally does not make its supported theory explicit, our ability to evaluate it at present is limited. For concreteness, we will outline what such a target theory might look like, a theory that we will refer to as $H_{3}$ and that addresses at least some of the most immediate inadequacies of the LLM Theory but is still more linguistically neutral than generative proposals. But we can of course have no certainty that this is what Wilcox et al. (2023), Mahowald et al. (2024), and others have in mind. We will therefore focus on methodological considerations that pertain to the evaluation of the target theory once it is provided and show how LLMs might bear on this evaluation. We show that, at least at present, LLMs lend no support to the Proxy View.
2.1 The LLM Theory and its inadequacy
As mentioned, Piantadosi (2023) argues for the LLM Theory of HLC. That is, he suggests that current LLMs, taken as real theories of HLC, are a better explanation of the known observations than anything coming out of the generative tradition. This claim is incorrect, for reasons discussed in detail in the literature and rehearsed briefly here. Following Katzir 2023 and Fox and Katzir 2024 we structure our brief discussion of the failure of the LLM Theory around three key issues: (a) competence vs. performance, (b) correctness vs. probability, and (c) learning and representations. [^1] See Kodner et al. 2023, Moro et al. 2023b, Rawski and Baumont 2023, and others for further critical discussion of the LLM Theory.
Competence vs. performance. Humans sometimes struggle with sentences that, given more time and focus, they ultimately accept. Familiar examples include center embedding, as in "The mouse that the cat that the dog chased bit died". Humans also sometimes initially accept sentences that, upon further thought, they reject. Familiar examples include variants of sentences such as the above but in which the verbs and the nouns are not balanced, as in "The mouse that the cat that the dog chased died". Other examples include so-called agreement attraction, as in "The keys to the cabinet is on the table".
The best explanation to these and other examples, going back to Yngve 1960 and Miller and Chomsky 1963, distinguishes linguistic competence from performance. The competence of English speakers, for example, licenses center-embedded structures regardless of depth, and it requires agreement between verbs and subjects regardless of intervening nouns. Performance factors get in the way and make center embedding and agreement across intervening nouns difficult and error prone. Among other things, the explanation in terms of a distinction between competence and performance accounts for the way center embedding can become easier or harder given various manipulations (tiredness, noise, parallel tasks, etc.).
Differently from common generative proposals, the design and behavior of current LLMs does not suggest a meaningful distinction between competence and performance of the kind discovered in humans: there is no reason to think that more (or less) time and working memory would make them change their behavior on center embedding or agreement. This leaves current LLMs with no good explanation for the increasing difficulty of sentences with increasing levels of center embedding or for the other observations that have been taken to support the distinction between competence and performance.
Correctness vs. likelihood. Humans recognize some sentences as correct but unlikely (one example again is center embedding as in the example above), and they recognize some other sentences as incorrect but likely (again, unbalanced variants of center embedding and agreement attraction, among other examples). LLMs provide straightforward information regarding their representation of likelihood. But it remains unclear whether they even have a distinct representation of correctness. [^2] A potentially important area of research concerns understanding the inner workings of LLMs, sometimes relating the discovered dynamics to linguistic patterns. See Lakretz et al. 2019 and Ravfogel et al. 2021, among others. It is conceivable that such investigations will eventually uncover a notion of correctness within LLMs.
[…]
Generative theories of HLC typically derive this inductive leap by assuming that children are born with a combination of representations and learning biases that favor the representation of constituency-based dependencies over dependencies that rely on linear order (in some approaches by preventing the latter from being represented in the first place). Current LLMs, on the other hand, seem to provide no basis for this inductive leap. In fact, recent work by Lan et al. (2022, 2024a) casts doubt on the ability of networks trained by current methods to even acquire the basic notion of constituency, let alone to explain the preference for constituency-based generalizations over linear-based ones. Not surprisingly, then, Yedetore et al. (2023) find evidence suggesting that at least some neural networks fail to acquire the correct constituency-based generalization from their training data. We will return to APS in section 3, where we will consider the possible role of LLMs in reasoning about the inductive leaps predicted by non-LLM theories of HLC.
Evidence about the representations and learning biases of humans also comes from the highly skewed distribution of linguistic patterns cross-linguistically. We can again consider constituency (following Fox and Katzir 2024), though we emphasize that a very wide range of cross-linguistic patterns have been discussed in the linguistic literature. We just mentioned that constituency is part of the best explanation of the competence of English speakers, noting its role in the formation of yes-no questions. Constituency is central to the explanation of many other aspects of English, including center-embedding and agreement mentioned earlier, wh-movement discussed in section 3 below, the computation of meaning, and more. But it seems clear that this is not just an accident of actual English: constituency is part of the best explanation of almost all languages that have been studied closely, which suggests that humans are born with representations and learning biases that favor constituency. Again, the representations and learning biases of current LLMs might not even be able to support the acquisition of constituency, let alone explain its cross-linguistic prevalence. We will return to cross-linguistic evidence in section 4, where we will consider the possible role of LLMs in reasoning about the typological patterns predicted by non-LLMs theories of HLC.
2.2 LLMs as proxies for good learners
But if the LLM Theory is discredicted, does this mean that LLMs are irrelevant to the scientific study of HLC? As mentioned above, the Proxy View explains why the answer is no: in principle LLMs can serve as proxies for theories of a third kind, theories that involve relatively linguistically-neutral representations and learning (and are different in this regard from generative theories) but that are not vulnerable to the most immediate arguments against the LLM Theory.
Here we consider two such uses of LLMs as proxies. Following Lan et al. 2024b we use LLMs as tools for reasoning about whether a relatively linguistically-neutral theory of HLC predicts a given inductive leap based on a given amount of linguistic exposure. And following Kallini et al. 2024 we use LLMs as tools for reasoning about whether a relatively linguistically-neutral theory of HLC derives a learning asymmetry that might be part of an explanation of a typological asymmetry.
At this point we should mention again an immediate obstacle to the proxy use of LLMs: while recent literature, including Wilcox et al. (2023) and Mahowald et al. (2024), suggests that the performance of LLMs supports linguistically-neutral theories of HLC, this literature does not provide much by way of specifics. Until such specifics are provided, expressions of hope about the undisclosed theories remain just that.
Still, we can outline some methodological considerations regarding the possible role of current LLMs in evaluating a linguistically-neutral theory of HLC once such a theory is made sufficiently explicit.
Let us illustrate with a concrete target theory, though we emphasize that this is for presentational purposes only and that we obviously cannot tell whether this target theory is at all close to what the literature supporting the Proxy View envisions. Consider then the following theory of HLC, which we will call $H_{3}$. According to $H_{3}$, the syntax is written and understood as Multiple Context-Free Grammar: any MCFG can be written, using the formalism of Seki et al. (1991), and no other grammar can be written. The semantics in $H_{3}$ is model-theoretic and is computed compositionally based on the syntax, while the syntax is blind to most of the semantics. Learning according to this theory follows the principle of Minimum Description Length (MDL; Rissanen, 1978). And processing relies on a parser with access to very little working memory and that can output not just parses (for grammatical inputs) but also probabilities.
$H_{3}$ is quite different from the LLM Theory and does not share the inadequacies of that theory reviewed above. For one thing, $H_{3}$ has a meaningful distinction between competence and performance: by relying on a limited amount of working memory, the parser will run into trouble with some of the same grammatically correct sentences that humans struggle with such as those with center embedding. $H_{3}$ also has a meaningful distinction between correct and probable. As to representations and learning, the combination of (a) MCFG as the representational framework for syntax and (b) MDL as the learning criterion will allow the learner to represent and learn constituency and perhaps also constituency-based dependencies. And the combination of those representations with compositional semantics can go a long way toward explaining the centrality of constituency across languages. As to modularity, this is baked into the assumptions mentioned earlier about the information available to the syntax. But while $H_{3}$ is very different from the LLM Theory, it is also more linguistically-neutral than common generative theories of HLC such as $H_{1}$. The representational system of $H_{1}$ would mostly likely include a specific preparation for grammars that involve linguistic patterns attested across langauges, such as constraints on movement commonly known as syntactic islands (Ross 1967), and including the phenomena we discuss in the following section. While $H_{3}$ makes it possible to state such grammars, it does not postulate anything specific to these patterns inside the theory, and does not provide further explicit biases beyond those present in a constituency-based syntax and model-theoretic semantics. It is therefore conceivable that the performance of LLMs might be informative to some extent about what can be learned and under what conditions under $H_{3}$, though adequately justifying the Proxy View for $H_{3}$ would require careful argumentation that goes beyond the present outline.
Criterion Generative Linguistics LLM Theory Proxy View
Linguistic Bias Strong Weak Intermediate
Competence vs. Performance $\checkmark$ $\times$ $\checkmark$
Grammaticality $\neq$ Likelihood $\checkmark$ $\times$ $\checkmark$
Explanation for learning assymetries $\checkmark$ $\times$ $\checkmark$
Concrete Instantiation $H_{1}$ $H_{2}$ $H_{3}$
Table 1: Comparison of the three theories outlined above on key issues, following Fox and Katzir 2024.
As mentioned, we have no way of telling whether $H_{3}$ is the theory of HLC that works such as Wilcox et al. 2023 and Mahowald et al. 2024 seem to have tried to promote on the basis of LLM performance. None of our methodological discussion below depends in any way on the specific properties of $H_{3}$: this theory can serve as a concrete example of a relatively linguistically-neutral theory of HLC if having such an example is helpful, but the reader should feel perfectly free to ignore it. It should also be clear that any serious attempt to promote $H_{3}$ or any other theory of HLC - as the best explanation would need to examine it in light of the whole body of work within theoretical linguists since the 1950s and that linguists have relied on to argue for theories of HLC that incorporate strong linguistic biases. Since our goal here is merely to illustrate the potential use of LLMs as proxies for linguistically-neutral theories, we will proceed as if this task has already been carried out.
3 Large Language Models and alignment with the stimulus
Consider then a theory of HLC - $H_{3}$ above if one so wishes, or something else -that is neutral with respect to some linguistic property $PROP$ for which generative theories include a strong linguistic bias. A potential argument against this relatively linguistically-neutral theory and in favor of a generative theory might take the form of an APS: if humans show knowledge of $PROP$ after exposure to a corpus that seems too impoverished to support the learning of $PROP$ by a linguistically-neutral learner, this will support the strongly biased generative theory. The problem is that it is generally very hard to determine what a given learner might acquire from a given corpus, even for a theory such as $H_{3}$ for which both the representations and the learning mechanism are explicitly provided.
LLMs might be able to help. We can train them on a suitable corpus and then probe their knowledge. Note that this is not entirely straightforward, for various reasons discussed in Lan et al. 2024b. Most immediately, while we might be able to directly inspect a grammar within a theory of HLC such as $H_{3}$ and ask whether it has knowledge of $PROP$, we cannot do the same with current LLMs, whose inner workings are extremely complex and remain mostly opaque, as discussed in Section 2.1. Instead of inspection, we might turn to behavioral measures such as acceptability judgments in minimal pairs where one member satisfies $PROP$ and is acceptable and a minimal variant violates it and is unacceptable, a methodology that has been used in Marvin and Linzen 2018, Hu et al. 2020 and Wilcox et al. 2019, among others. But while the theory of HLC under consideration might offer such judgments, current LLMs provide likelihood assessments, which are an entirely different thing, as mentioned above. However, if we focus on areas where acceptability and likelihood are reasonably well aligned, we might be justified in looking at whether the LLM assigns a higher probability to the $PROP$-satisfying member of the pair than to the $PROP$-violating one. If this probabilistic preference is sufficiently strong, we might consider the fact that the LLM has acquired this preference as reason to think that the learner under our target theory of HLC would acquire the actual $PROP$. This might be so even if the probabilistic preference of the LLM reflects a flawed approximation of $PROP$ rather than knowledge of $PROP$ in any meaningful sense.
Following Wilcox et al. and Lan et al., our success criterion considers the model successful if and only if it assigns a higher probability to the grammatical member of the minimal pair. Specifically, we evaluate probabilities on a critical region token whose preceding context is the same in both the grammatical and the ungrammatical sentence (see Table 3, where critical regions are indicated with underscores). Note that this is an extremely lenient notion of success: the model is considered successful even if the preference for the grammatical member of the minimal pair is extremely small, and even if ungrammatical sentences are considered more likely than some (less related) grammatical sentences.
3.1 Methods
Model Train Dataset Size Human Equivalent
CHILDES LSTM, Transformer 8.6 million tokens 10 months
BabyLM 10M 10 million tokens 1 year
Wikipedia Transformer 90 million tokens 8 years
BabyLM 100M 100 million tokens 9 years
bert-based-uncased $\approx$ 3.5 billion tokens 320 years
GPT2 $\approx$ 8 billion tokens 730 years
llama3.2-3b $\approx$ 9 trillion tokens 821,250 years
Table 2: Training data size of the eight language models considered here, and the human linguistic experience equivalent to these data sizes, following Hart and Risley.
Let us illustrate this kind of reasoning with concrete examples. To do so, we use an empirical setup that consists of eight models of different architectures and training schemes (see Section A). [^4] All experimental material can be found in https://osf.io/5zh6q/. The models were trained on datasets that roughly correspond to different timeframes of linguistic experience, ranging from 10 months of human experience (CHILDES) through eight years of linguistic experience (English Wikipedia subset) to 320,730 or even 821,250 years of linguistic experience (BERT base uncased, GPT-2 and Llama 3.2-3b respectively). [^5] More concretely, we use two models from Yedetore et al., an LSTM and a Transformer, that were trained on the CHILDES corpus of child-directed speech; a Transformer we trained on a subset of English Wikipedia, for which we used one of the large Transformer architectures used in Yedetore et al.; the base-strict and base-strict-small versions of the BabyLM LLMs from Warstadt et al. (2023a), trained on the 100M and 10M BabyLM datasets correspondingly; the BERT base uncased version of BERT (Devlin et al. 2018a), trained on the BooksCorpus and English Wikipedia datasets; OpenAI’s GPT-2 (Radford et al. 2019a); Meta’s Llama 3.2-3b LLAMA model (released September 2024). See Section A for further technical detail. The models were probed on their knowledge of three linguistic phenomena using the success criterion introduced above (see Table 3). The linguistic test cases presented in the sections below show that all models, despite being trained on multitudes of data, perform around chance level and worse on the lenient success criterion. This result is not optimistic for LLMs as a proxy for an adequate theory of HLC.
Phenomenon Example (Grammatical/*Ungrammatical)
Across-the-board movement (ATB) Which boy did you say that Kim hated and Mary loved yesterday?
* Which boy did you say that Kim hated and Mary loved Ann yesterday?
Parasitic gaps (PG) I know who John’s talking to is going to annoy soon.
* I know who John’s talking to is going to annoy you soon.
That-trace effects (TTE) Who did you say that Sue loves?
Who did you say loves Sue?
Who did you say Sue loves?
* Who did you say that loves Sue?
Table 3: Examples of linguistic phenomena with grammatical and ungrammatical sentences. Model probabilities are compared on the underscored critical regions.
3.2 Test Case: Across-the-board movement and parasitic gaps
[…]
The common analysis of wh-movement in generative theories involves a constituent that is attached in one position and is then re-attached higher up in the structure. In (3.2), the constituent [which book] is attached within the embedded clause as a sister to Mary and is then re-attached at or near the root of the matrix clause, which among other things leads to it being pronounced in the beginning of the sentence.
Some work within generative linguistics (Pearl and Sprouse 2013, Phillips 2011) suggests that a relatively linguistically-neutral theory cannot account for the full knowledge of wh-movement given the kind of data that children are exposed to, an instance of the APS. This is not a consensus view, however, and some recent work such as Wilcox et al. suggests that LLMs undermine this APS and that these models show that a linguistically neutral theory would, in fact, acquire a full knowledge of wh-movement from a developmentally-realistic corpus. Following Lan et al. let us show how LLMs (used as proxies for an undisclosed linguistically-neutral theory of HLC) might bear on the APS under consideration and why they currently do nothing to undermine it.
Before proceeding, let us reiterate that it is hard to tell whether LLMs are good proxies for the relatively linguistically-neutral theory of HLC that is envisioned in works such as Wilcox et al. (2023) in the absence of at least some specifics about that theory. But if they are, and if we focus on cases where correctness and likelihood are reasonably well-aligned, we could reason that if the LLM overwhelmingly assigns much higher probability to the correct member of each minimal pair, this indicates that there was sufficient information in the training data for the target linguistically-neutral theory to have acquired the relevant knowledge. And if the LLM does not perform as well, this suggests that the target theory would also struggle. The former possibility weakens the APS, and the latter strengthens it.
[…]
We test four models from Table 2 on their knowledge of across-the-board movement and parasitic gaps using the minimal pair method. We compare sentences that both have an upstream filler - like which book in Example 3.2, or who in Example 3.2 - and differ in whether they are gapped accordingly. [^6] The paradigm sentences in Lan et al. (2024b) were generated by template for ATB and PG. Here we unify the templates for both phenomena so that they use the same lexical choices where relevant (e.g., proper names), and by adding more lexical choices. From each template, represented by a context-free grammar, we sampled 10,000 sentence pairs and ran them through each model to get the relevant probability values (see Appendix B). As is shown in Figure 1, all prefer the ungrammatical over the grammatical members of the minimal pairs in the vast majority of cases. This failure suggests that the linguistically-neutral theory for which the LLMs are taken to be useful proxies would not acquire wh-movement from a developmentally-realistic corpus.
Figure 1: Model accuracy on ATB and PG datasets averaged over five experiment seeds. Accuracy is measured as the ratio of cases where the model assigns a higher probability to the grammatical sentence continuation.
3.3 Test Case: that-trace effects
Another property of wh-movement in English is that when the moved wh-phrase is a subject, the clause from which it moves cannot have an overt complementizer such as ‘that’:
[…]
This asymmetry, known as that-trace effects (TTE), has been thoroughly discussed in theoretical linguistics (see Perlmutter 1968, Chomsky and Lasnik 1977, and Phillips 2013a, among others). It has been shown that extraction of any kind from clauses with that is vanishingly rare in adult-directed corpora, and even rarer in child-directed speech Phillips (2013a) [^7] Out of 11,308 analyzed utterances from child-directed speech, there were only two instances of non-subject extraction with that and zero instances of subject extraction with that, contrasting with 159 instances of object extraction without that, and 13 instances of subject extraction without that.. This distribution is claimed (Chomsky and Lasnik 1977, Phillips 2013b) to be insufficient to warrant acquisition of the paradigm above. We set aside the important theoretical discussion of what bias present in the architecture or representations of $H_{1}$ underlies the TTE according to generative linguistics. [^8] See Sobin (1987), Lohndal (2009), Bošković (2016) and Pesetsky (2017), among others, for discussion. Instead, we focus on its status as an APS, which warrants evaluating the performance of the linguistically-neutral LLM contender on it.
Similarly to the previous case, we test eight LLM models (see Table 2). The four-way asymmetry warrants two criteria that constitute knowledge of the TTE: first, given an overt complementizer upstream, the model should prefer an overt subject to a gap in the embedded clause. This preference should translate into probabilities so that $P(loves)$ in 3.3 should be smaller than $P(Sue)$ in 3.3. Second, given a gapped sentence downstream ("…       loves Sue"), the model should prefer a prefix without an overt complementizer. That is, $P(loves)$ in 3.3 is expected to be smaller than $P(loves)$ in 3.3. This way, we measure the dispreference for an overt complementizer in the presence of a moved subject along both axes - that of the complementizer and that of the gap.
Results for both success criteria are presented in Figure 2. Under both criteria, the performance of the models hovers around chance. While it is unclear how to interpret these results, we see nothing in the performance of the models that suggests that the target hypothesis will account for the learning of that-trace effects from similar training corpora.
Figure 2: Model accuracy values for the $P(+that,+trace)<P(+that,-trace)$ criterion (top) and $P(+that,+trace)<P(-that,+trace)$ criterion (bottom), over samples of 10,000 sentence pairs from the TTE test set. Results are averaged over five seeds. The dark plot represents model training sizes.
4 Large Language Models and cross-linguistic evidence
Consider again the comparison of a strongly linguistically-biased generative theory and a more linguistically-neutral alternative. Suppose that the strong linguistic biases of the generative theory can be shown to help derive a cross-linguistic pattern, perhaps by making it easier to represent and learn the languages that adhere to the pattern than those that do not. In the linguistically-neutral theory it can be much harder to see directly whether the attested patterns are predicted to be easier to learn than the unattested ones. One might try to evaluate this matter empirically, by simulating learning of languages of the attested and the unattested kind given the linguistically-neutral theory, but this can be hard to do. Consequently, it can be hard to see whether the theory can explain the cross-linguistic pattern under consideration and whether it can compete with the generative theory. Again, however, if an LLM is a good proxy for the theory, the problem is avoided to some extent. One can examine the ease with which the LLM learns to approximate both languages that follow the pattern and languages that violate it. If the LLM approximates the former more easily than the latter, this can suggest that the target theory might be able to account for the cross-linguistic pattern (subject to various further assumptions about how relative ease of learning derives the typology, and to an adequate definition of how ease-of-learning via LLM is to be measured). If, on the other hand, some of the pattern-violating languages are easier to learn than some of the pattern-satisfying ones, the linguistically-neutral theory would need further mechanisms to derive the typology.
We should emphasize that the inferences from this argument are particularly weak. On the one hand, even if the linguistically-neutral theory derives a learning asymmetry between pattern-satisfying and pattern-defying languages but the difference is very small, this might not suffice to account for a robust typological pattern. On the other hand, even if the linguistically-neutral theory does not derive a learning asymmetry, it might be able to account for it in terms of other factors, such as communicative pressure. In order to overcome these limitations one could try to situate the learning component within a model of cultural evolution, which allows for the amplification of small asymmetries over generations and also factors in communicative pressures and other considerations. We believe that such a model is the proper context for the use ease of learning by LLMs to evaluate typological asymmetries, but following Kallini et al. (2024), whose setup we use here, we leave such a model for future work and only assess currently available models. [^9] Kallini et al. (2024) contextualize their work differently. We will not attempt to determine whether the present work is in line with their goals.
4.1 Ease of learning by LLM as a theory of HLC?
Kallini et al. (2024) compare GPT-2 when trained on an English corpus and when trained on corpora that are derived from the original corpus through various perturbations, each corresponding to an unattested language. They find that the original English corpus was easier for GPT-2 than the derived corpora of unattested languages when ease-of-learning is measured through test set perplexities through time (see Figure 3). In doing so, Kallini et al. (2024) make use of LLMs as proxies in the context of explaining the cross-linguistic skew that rules out the unattested perturbations of English, even though this theory is never made explicit in the paper. If a model like GPT-2 is indeed a good proxy for this contender, then we may say that what Kallini et al. (2024) have shown to hold for English should also hold cross-linguistically, as presented in the introduction: all unattested languages should be harder to learn than attested ones.
In the following subsections we consider a few languages for which the evaluation is simple given current resources and test them with the perturbations used by Kallini et al. (2024) and one additional perturbation, listed in Table 4. For each perturbation, we create perturbed versions and baseline versions for each of the four languages - English, Italian, Russian and Hebrew, based on datasets from Gulordava et al. (2018). We then train a transformer LLM based on the architecture and training scheme from Yedetore et al. (2023), and consider the validation set perplexities as training progresses. [^10] The model was trained for 48 hours and validation perplexity was computed every 200 training batches. The 48 hours of training amounted to roughly 1.5 epochs per dataset, with batch size 10, such that each epoch amounted to an average of 181,552 batches, see Appendix B for further details. We find we find several different languages of an unattested kind that are easier for the LLM to approximate than a corresponding attested language, as measured by the perplexity metric used by Kallini et al. (2024). From the perspective of the Proxy View, this means that the target, relatively linguistically-neutral theory fails to receive support from the LLM: if the target hypothesis can explain the typological pattern, this is not detected by the LLM experiment.
Perturbation Attested / Perturbed
partial-reverse Colorless0 green1 <rev>2 ideas3 sleep4 furiously5.
Colorless0 green1 <rev>2 furiously5 sleep4 ideas3.
full-reverse Colorless0 green1 <rev>2 ideas3 sleep4 furiously5.
Furiously5 sleep4 ideas3 <rev>2 green1 colorless0.
switch-indices Colorless0 green1 ideas2 sleep3 furiously4.
Ideas2 green1 colorless0 sleep3 furiously4.
token-hop They0 were1 sleeping2 v3 next4 to5 the6 colorless7 green8 ideas9.
They0 were1 sleeping2 next3 to4 the5 v6 colorless7 green8 ideas9.
Table 4: Perturbation test cases. Ease-of-learning is evaluated for the attested and perturbed versions of English, Italian, Hebrew and Russian.
4.2 Test case: partial-reverse
We demonstrate the ease-of-learning method in our first test case - a cross-linguistic reproduction of the partial-reverse perturbation from Mitchell and Bowers 2020 and Kallini et al. 2024. Each sentence is reversed starting from a randomly chosen index in the input sentence, in which a special marker token <rev> is inserted:
[…]
The target, relatively linguistically-neutral theory in this case might be a version of a generative theory in which there is a parameter for reverse. Any grammar that is statable within the theory now comes in two variants. If the parameter is set to 0, the grammar works as in the original generative theory. But if the parameter is set to 1, every derivation ends with reversing the output starting from a random point. This may or may not be the target theory that Kallini et al. have in mind; other target theories are of course imaginable, and Kallini et al. do not comment explicitly about this. We note that, like Kallini et al.’s other perturbations, reversal seems quite far away from the typological questions that generative linguistics have usually concerned themselves with. Here we stay with Kallini et al.’s perturbations and setup simply in order to illustrate the relevant methodological considerations, but a proper exploration of the Proxy View would need to look at those cross-linguistic patterns that have informed linguistic research.
[…]
We find that validation perplexities follow an attested-unattested divide: the partial-reverse perturbed versions of English, Italian and Russian are harder to learn than their attested counterparts (see Fig. 3). [^11] The partial-reverse languages are compared to a modified baseline of English in which the marker <rev> is inserted randomly without reversing, to control for the effect for additional textual material in the sentence on perplexities. Since we use the same seed for both perturbations, the markers are inserted in the same indices (see example 4.2).
This, as discussed above, could be part of an argument that linguistic biases are not needed to rule out partial-reverse languages, but crucially only if there were a serious contender for the best explanation that allows for such languages in the first place, and could be compared theoretically with the restrictive generative account. However, other perturbations paint a different picture under the cross-lingustic view.
4.3 Test case: full-reverse
We now consider the perturbation $\Pi=$ full-reverse from Kallini et al. 2024, where each sentence is reversed in its entirety and a special marker token is randomly inserted:
[…]
We find that the perturbed versions of Italian, Russian and Hebrew are projected as easier than attested English (see Fig. 4), even though the perturbation is considered humanly impossible. In this case, ease of learning according to Kallini et al. (2024)’s perplexity metric does not provide support for the idea that the target theory can explain away the typological asymmetry through asymmetries of ease of learning. As mentioned above, this conclusion is weak. It is possible that the target theory is the best explanation and that the typology is explained with the help of factors other than ease of learning. It is also very possible that Kallini et al.’s perplexity metric is a poor way to evaluate ease of approximation by the LLM. All we can conclude at present is that proponents of the Proxy View have work to do in light of the LLM’s performance in this case, and that previous results only provide a limited picture.
Figure 4: Validation perplexity during training for attested (baseline) and full-reverse versions of English, Russian, Italian, and Hebrew.
4.4 Test case: switch-indices
We turn to another asymmetry that the Proxy View does not at present explain. Consider $\Pi=$ switch-indices, where the tokens at index 0 and index 2 in every sentence are switched:
[…]
One can again imagine various target theories that allow for languages like English as well as their switch-index variants. And again the question would be whether such a target theory can account for the typological observation that such switch-index variants are unattested.
If the target theory can account for the typological asymmetry, this is again not detected through the perplexity curve of the LLM proxy: each language is wrongly projected as harder than its perturbed version, and the perplexity delta between each two attested languages is significantly larger (see Fig. 5). In all cases, there is an unattested language that is projected as easier than an attested language.
Figure 5: Validation perplexity during training for the attested (baseline) and switch-indices versions of English, Italian, Hebrew and Russian.
4.5 Test case: token-hop, no-hop
[…]
Since additional textual material is inserted, in a way that might confound perplexities over time, we compare the perplexities achieved by our model on token-hop datasets to no-hop datasets, in which the marker token is inserted exactly after each verb. This achieves control for the effect of the marker token itself, such that difference in perplexities can be attributed solely to the effect of the token-counting generalization. The no-hop version adheres to attested generalizations, as it can be thought of as post-verbal clitic that marks for parts of speech. Learning token-hop on the other hand crucially requires counting tokens, an ability that is particularly non-humanlike, both because human languages have no atomic concept that corresponds to tokens, and because counting rules violate the strictly hierarchal nature of linguistic generalizations across languages. Although the distance of the token from its associated verb has an effect on perplexity, we find that the no-hop version of English is still harder to learn than token-hop Italian and Russian (see Fig. 6). Unperturbed English datasets are consistently projected as harder than token-hop perturbed versions of Italian and Russian, despite a clear typological skew against counting rules. The same point is made with the validation perplexities of token-hop Italian and no-hop Russian patterning very similarly. The best explanation of linguistic typology should predict a clear divide between unattested languages that crucially require counting linguistic elements and attested languages. Ease-of-learning via LLM does not explain this clear typological skew.
Figure 6: Validation perplexity during training for the no-hop (baseline) and token-hop versions of English, Italian, and Russian.
5 Conclusion
LLMs have been presented as a challenge to more traditional approaches to the study of HLC, and in particular to generative linguistics. A blunt version of this idea is the LLM Theory, promoted by Piantadosi (2023), which maintains that current LLMs are themselves good theories of HLC. Current LLMs can, of course, be viewed as theories of HLC, but this treatment does these models no favors, as the literature was quick to note.
A more cautious approach is the Proxy View, promoted by Wilcox et al. (2023) and Mahowald et al. (2024), which maintains that current LLMs are good proxies - as far as various behavioral measures are concerned - for good theories of HLC that are more linguistically-neutral than generative theories. No commitment to shared essentials between LLMs and the relevant good theories is implied. The Proxy View is of course perfectly coherent, but differently from the LLM Theory it is too vague about the theories of HLC that it champions to be of much direct use within empirical science. Empirical science is based on inference to the best explanation and proceeds through competition between reasonably well-understood theories, and saying just that one’s favorite theory learns roughly as well as an LLM is insufficient for this purpose. One would like to see enough detail at least for a proper comparison of the relevant theory to generative linguistics in light of the discoveries and considerations in the literature of the past 70 years or so. More than anything else, the present paper is a plea for supporters of the Proxy View to provide this kind of detail and evaluation for the theory they have in mind.
Still, already at this stage we can outline an empirical evaluation of the Proxy View on various kinds of test cases, and here we focused on two: alignment with the stimulus and cross-linguistic variation. The results were not encouraging for the Proxy View. The LLMs failed to even approximate key pieces of knowledge that humans have, even when the models were trained on much larger corpora than children are exposed to. The LLMs also had an easier time approximating various typologically-unattested languages than actual human languages, at least with respect to a success criterion used in Kallini et al. (2024).
Appendix A Section 3 - Appendix
A.1 General Notes
In all experiments, probabilities are evaluated on the critical region in boldface (as in the example sentences in Tables A.2, A.2, A.2). For models listed as unidirectional, the probability value takes into consideration the left context of the evaluated token, whereas for bidirectional models, the probability value factors in both the left and the right context. The bidirectional models we consider were pretrained with masked language modeling (fill-mask).
In all generated sentences, the critical region unambiguously disambiguates the sentence as an instance of the phenomenon under consideration (there is no other grammatical reading). To promote lexical variation in target regions, possible target tokens in the CFGs were chosen at random from a POS-tagged English Wikipedia dump. We filtered out potential target tokens that were not part of the vocabulary of one the models considered in our setup, which are detailed in Table 5.
Model Name Training Size Directionality Source
CHILDES LSTM 8.6M Uni Yedetore et al. 2023
CHILDES Transformer 20M Uni Yedetore et al. 2023
BabyLM 10M (strict-small) 10M Bi Warstadt et al. 2023b
Wikipedia Transformer 15M Uni Trained by authors
BabyLM 100M 100M Bi Warstadt et al. 2023b
BERT Base Uncased $\approx$3.5B Bi Devlin et al. 2018b
GPT2 $\approx$8B Uni Radford et al. 2019b
Llama 3.2-3b $\approx$9T Uni Touvron et al. 2023
Table 5: Models used in the experiments in Section 3.2.
A.2 ATB and PG
For the ATB and PG experiments, we replicate the experimental setup from Lan et al. 2024b with a wider variety of test sentences to increase robustness. We also unify the templates for both phenomena so that they use the same lexical choices where relevant (e.g., proper names). From each template, represented by a context free grammar, we generated tuples of sentences that correspond to the (+filler,+gap), (+filler,-gap) conditions (see CFGs and example sentences in Tables A.2 and A.2). We then evaluated the success criterion $P(+filler,-gap)<P(+filler,+gap)$ on a sample of 10,000 tuples for each phenomenon. For ATB, sentence structure was crucially changed such that the two conjuncts do not agree in number, to prevent a problematic non-ATB reading present in some of the Lan et al. test sentences. Underlined words alternate according to the ±$\pm filler$ condition; words in bold mark the position where the ±$\pm gap$ condition becomes evident and probability is measured. We evaluate the performance of four models from Table 5. We do not consider bidirectional models for the PG and ATB experiments because they are highly sensitive to whether the token is the last token in the sentence, which in the ATB and PG case is one of the minimal differences between the conditions we consider.
ATB Grammar
$S\rightarrow\langle{PREAMBLE}\rangle\ \langle{\pm F}\rangle\ \langle{LINK}\rangle\ \langle{\pm G}\rangle$
$\langle PREAMBLE\rangle\rightarrow\ \textit{I know}$
$\langle{+F}\rangle\ \rightarrow\ \underline{\textit{{which}}\ \langle{PLNOUN}\rangle\ }\langle{VP1}\rangle\ \ \langle{ADV1}\rangle$
[…]
$\langle{-G}\rangle\ \rightarrow\langle{LINK}\rangle\ \ \langle{VP2}\rangle\ \ \langle{\textbf{OBJ}}\rangle\ \ \langle{ADV2}\rangle$
[…]
$\langle{VP1}\rangle\ \rightarrow\langle{VP1\_SIMPLE}\rangle\ \ |\ \langle{VP1\_COMPLEX}\rangle$
[…]
$\langle{VP2}\rangle\ \rightarrow\langle{VP2\_SIMPLE}\rangle\ \ |\ \langle{VP2\_COMPLEX}\rangle$
[…]
$\cdots$
$\Rightarrow$ I know which boys John met recently and is going to hug soon. (+filler,+gap)
$\Rightarrow$ *I know which boys John met recently and is going to hug you soon. (+filler,-gap)
Table 6: Excerpt from the context-free grammar used to generate Across-the-Board sentences for the experiments in Section 3.
Table 7: Excerpt from the context-free grammar used to generate Parasitic Gap sentences for the experiments in Section 3.
A.3 TTE
To create test sets, we generate 1,178,496 quadruplets, out of which we randomly sample 10,000 quadruplets per experiment. Paradigms are varied lexically in several ways: we first create paradigms that differ structurally from each other, to control for the possibility that our results are an artifact of the specific sentence structure introduced in the example sentences in Table A.2. We also introduce lexical variation to all non-target elements, such as the main subject, the embedding verb, the wh-element and the verb inside the wh-clause. Most importantly, lexical choices for target verbs and nouns are varied using 30 target nouns and 30 target verbs. We perform five inference experiments on five such samples and consider average accuracy over all experiments. Quadruplets are evaluated with the two success criteria introduced in Section 3.3.
Table 8: Excerpt from the context-free grammar used to generate That-Trace-Effect sentences for the experiments in Section 3.
Appendix B Section 4 - Appendix
B.1 Dataset Creation
Our baseline datasets are based on Wikipedia dumps in English, Italian, Hebrew and Russian from Gulordava et al. (2018), which extracted 90M token subsets for each language, and switched all tokens that do not belong to the top 50K most frequent tokens into $\langle{unk}\rangle$ tokens (see Gulordava et al. 2018). For each baseline dataset we create perturbed train, test and validation datasets (see Table 4), similarly to Kallini et al. 2024, such that each perturbed dataset has the same number of tokens as its corresponding baseline dataset.
[…]
  1. 1.
    "en_core_web_sm", accuracy of POS tagger: 0.97
  2. 2.
    "it_core_news_sm", accuracy of POS tagger: 0.97
  3. 3.
    "ru_core_news_sm", accuracy of POS tagger: 0.99
We chose to replace Kallini et al. (2024)’s 3rd person agreement token-hop with a POS marker because of the lower accuracy of morphological taggers on languages that are not English and cross-linguistic differences in verb morphology.
B.2 Model Training and Evaluation
We use the optimal Transformer architecture chosen in a hyperparameter search conducted by Yedetore et al. 2023, which has four layers, a hidden and embedding size of 800, a batch size of 10, a dropout rate of 0.2 and a learning rate of 5. Every 200 batches, we evaluate the average perplexity per token of the current trained model on a held out validation set. We set a 48-hour training threshold, during which each model processed a different number of batches. For consistency in comparison graphs, we use the smallest batch count achieved and compare all models up to that point. Table 9 details the number of training batches achieved by the Transformer model on each perturbed dataset.
Language Perturbation Size of Validation Set (batches) Training batches
English no-perturb 138430 137600
full-reverse 143518 117400
partial-reverse 143517 131800
switch-indices 166116 205516
no-hop 181502 189102
token-hop 181502 142200
Italian no-perturb 138589 137200
full-reverse 143840 154200
partial-reverse 143506 152000
switch-indices 166307 213507
no-hop 181552 146400
token-hop 181552 191752
Russian no-perturb 138602 136200
full-reverse 145993 153600
partial-reverse 145984 125400
switch-indices 166323 213323
no-hop 178771 148800
token-hop 178771 192571
Hebrew no-perturb 137773 138000
full-reverse 142884 154800
partial-reverse 171460 203860
switch-indices 165328 210128
no-hop - -
token-hop - -
Table 9: Number of training batches achieved in 48 hours of training on our perturbed datasets from Section 4.